OpenStudy (sawq):

Urgent! Will fan and medal!! x + y = 27 4x + 1y = 45 Whats the solution to the equations and how did u get it? I know about substitution and elimination but its not working out... Help...??

pooja195 (pooja195):

What would you like to use?

pooja195 (pooja195):

I suggest elmination

OpenStudy (sawq):

Ok elimination then

pooja195 (pooja195):

Keep this in mind: OPPOSITES CANCEL whats the opposite of 4 ?

OpenStudy (sawq):

-4

OpenStudy (usukidoll):

I second elimination method.. works best.

pooja195 (pooja195):

\[-4(-x + y = 27)\] distribute

OpenStudy (sawq):

4x + -4y = 27 ?

OpenStudy (usukidoll):

close but you forgot to distribute the -4 to the 27 what's -4 x 27?

pooja195 (pooja195):

i accidently slipped in the -

pooja195 (pooja195):

-4(x + y = 27) -4 * x = ? -4 * y = ? -4 * 27= ?

OpenStudy (sawq):

-108

OpenStudy (usukidoll):

we are distributing the -4 all over the first equation.

OpenStudy (usukidoll):

and yes -4 x 27 is -108 :)

OpenStudy (sawq):

-4x -4y -108

OpenStudy (usukidoll):

correct so your new first equation is -4x-4y=-108

pooja195 (pooja195):

ill let doll finish :)

OpenStudy (usukidoll):

so now we have -4x-4y=-108 4x + 1y = 45

OpenStudy (usukidoll):

what do we need to do to get rid of the x?

OpenStudy (sawq):

lol @pooja195 thanks for ur help

OpenStudy (sawq):

inverse? @UsukiDoll

OpenStudy (usukidoll):

umm .. not quite.. our goal is to eliminate a variable.

OpenStudy (usukidoll):

-4x-4y=-108 4x + 1y = 45 so what do we need to do to get rid of this x?? we have -4x and 4x... maybe adding these equations together might work

OpenStudy (sawq):

ooh right so it would be: x - 5y = -63?

OpenStudy (usukidoll):

hmm .... only the right hand side of the equation is correct

OpenStudy (usukidoll):

ok... let's think finance for a bit you bought pens for $4, you paid $4 -4x+4x = ?

OpenStudy (sawq):

0x or just x

OpenStudy (usukidoll):

0x is fine... but 0 times x => 0

OpenStudy (sawq):

ohhhh

OpenStudy (usukidoll):

ok so we got rid of the x... now for the y we have -4y+1 it's like you are trying to buy pens for $4, but all you have is a dollar. How much more money do you need?

OpenStudy (usukidoll):

oops I meant -4y+1y

OpenStudy (sawq):

3 dollars

OpenStudy (sawq):

so -3y

OpenStudy (usukidoll):

yes! so -4y+y = -3y

OpenStudy (usukidoll):

so now we have -3y=-63

OpenStudy (usukidoll):

all we need to do is isolate y ... how do we do that?

OpenStudy (usukidoll):

in other words... we need y by itself.. no numbers no variables..

OpenStudy (sawq):

divide -63/-3=21

OpenStudy (usukidoll):

mhm so y = 21

OpenStudy (usukidoll):

the second part is finding x.. luckily it's easier.. we just need to subsitute y = 21 in one of the original equations. I highly suggest the easier ones

OpenStudy (usukidoll):

x + y = 27

OpenStudy (sawq):

ok: x + 21 = 27 x= 6

OpenStudy (usukidoll):

yes

OpenStudy (usukidoll):

so x =6 and y =21, but do these values work for the other equation. I know it does for the first equation because x+y=27 6+21=27 27=27

OpenStudy (usukidoll):

4x + 1y = 45 so plug in x = 6 and y =21

OpenStudy (sawq):

4*6=24 1*21=21 21+24=45!!!

OpenStudy (usukidoll):

yes 21+24=45 45=45 so we have found our x and y values.

OpenStudy (usukidoll):

alternatively you can choose to eliminate the y first and solve for x

OpenStudy (usukidoll):

x + y = 27 4x + 1y = 45 -x-y=-27 4x+y=45 3x = 18 x=6 6 + y = 27 -6 -6 y=21

OpenStudy (usukidoll):

^_^ I got the same values by eliminating the y first and my steps were shorter too.

OpenStudy (sawq):

Thanks a million @UsukiDoll you are the best!!!!! you really helped me

OpenStudy (usukidoll):

I've noticed that through inspection I had two one y's so I figured if I make one of the y's negative I can solve this quickly.

OpenStudy (usukidoll):

:D you're welcome

OpenStudy (sawq):

gnite tysm